OpenStudy (anonymous):

How do I solve integral from 0 --> 2pi 1/(1+cosx)dx @Hero @Directrix @Nurali @Luigi0210 @agent0smith @dan815 @primeralph @Compassionate @shamil98 @nincompoop @ranga @Ashleyisakitty @Isaiah.Feynman @wolfe8 @ehuman @RANE @wolf1728 @RolyPoly @UsukiDoll @amoodarya @missMob @Rina.r @Mr.ClayLordMath @Kristen17 @Victoria_Davis_2018 @AccidentalAiChan @khadeeja @doc.brown @RyGuy @khadeeja @adrynicoleb @ComeAlongP0nd @celecity @JMark @JessicaBBY @LanguageEnthusiast @averixx @arilove1d @Stormswan

OpenStudy (shamil98):

You know, you don't have to resort to mass tagging with your question each time. If there are people capable and available of assisting you they will probably do so..

OpenStudy (anonymous):

Excellent! So what's the answer then, @shamil98?

OpenStudy (anonymous):

We don't just give away answers

OpenStudy (shamil98):

\[\huge \int\limits_{0}^{2\pi} \frac{ 1 }{1 +\cos x} dx\] You are asking "how" to solve it? right?

OpenStudy (shamil98):

Use u-substitution.

OpenStudy (anonymous):

Are you guys dumb? U-Substitution won't work, and @Victoria_Davis_2018, since you don't give answers why don't you show me how to do it? If you don't know how, don't spam on my page, thanks.

OpenStudy (usukidoll):

NOOOOOOOOO don't just give out hints...ask him what's up doc?

OpenStudy (anonymous):

I wouldn't think that a future math professor would appreciate handing out answers, it impedes the learning process. Just sayin'

OpenStudy (usukidoll):

he wants to integrate foooooooolz

OpenStudy (usukidoll):

I'm not giving him no bailout or some shiz.

OpenStudy (amoodarya):

careful about vertical retricemptote

OpenStudy (usukidoll):

wolfram it and learn it from there ^^

OpenStudy (usukidoll):

definite integral so yeahhhhh.

OpenStudy (amoodarya):

may be in complex !

OpenStudy (anonymous):

@ehuman I've done all the work, by giving me the answer it makes me learn better then wasting my time showing me YOUR process. It's probably inefficient anyways.

OpenStudy (anonymous):

show your work and your answer then. easier to compare YOUR efficient process to our inadequate ones.

OpenStudy (anonymous):

You would just say "that's right" without even looking at it, then why would I waste my time posting on here then? That's inefficient.

OpenStudy (usukidoll):

omg drama ahead

OpenStudy (stormswan):

kitty got a temper......

OpenStudy (anonymous):

up tight much? you may have some maturing to do before the grave responsibility of teaching students is thrust upon you.

OpenStudy (usukidoll):

XD

OpenStudy (anonymous):

moving along, nothing to see here.

OpenStudy (anonymous):

it's not really a fight if one is closing his eyes and swinging at air is it?

OpenStudy (anonymous):

wow never saw that many that trying answer one question

OpenStudy (anonymous):

I really need help with a question im working on can someone please help ?

OpenStudy (anonymous):

when i call some one no one comes -.-

OpenStudy (doc.brown):

Please don't include me in mass tagging.

OpenStudy (stormswan):

same here.....^

OpenStudy (usukidoll):

SAME TOO! LIKE @*(#$&#@*(#@ STOP N STUFF MAN! FOR ALL WE KNOW IT CAN'T BE INTEGRATED OR SOMETHING OR NOBODY KNOWS

OpenStudy (anonymous):

Yeah, I got called here for the right to have a tantrum thrown at me. Merry Christmas all. You are here, it's late. It shows you are committed to learning or helping others learn, or you didn't finish your homework in time and need fast answers.

OpenStudy (stormswan):

...but everyone is on Christmas break...

OpenStudy (anonymous):

cosx = 2cos^2 (x/2) - 1

OpenStudy (anonymous):

Hope this will help u...:)

OpenStudy (primeralph):